Raffle tickets for a trip to Miami are assigned three digit numbers using the digits 0-9. Find the probability that the number on the winning ticket is either even or less than 300?
Answer #1

Total no of tickets are = 10*10*10 = 1000. (Assuming 005 is also a 3 digit no.)

Number of numbers less than 300 = 300 (0 to 299)

Number of even numbers from 300 to 999 = 350

Number of numbers either even or less than 300 = 350+300 = 650

Probability = 650/1000 = 0.65